OpenStudy (anonymous):

What are the x-intercepts for the graph of y=3(x-1)(x+6)? A. (-1,0) and (6,0) B. (1,0) and (-6,0) C. (0,-1) and (0,6) D. (0,1) and (0,-6) ***Idk how to solve.. please explain? thanks ;)

OpenStudy (anonymous):

x intercepts are zeros of the functions. Now, for which values of x will y be 0?

OpenStudy (anonymous):

idk :(

OpenStudy (anonymous):

umm -1 and 6?? idk :/

OpenStudy (anonymous):

no, if you inserted -1, how much would you get? 6 is correct

OpenStudy (anonymous):

uh no, sorry

OpenStudy (anonymous):

-6 would be correct

OpenStudy (anonymous):

-1 as x?

OpenStudy (anonymous):

yes, how much would y be if x=-1

OpenStudy (anonymous):

\[0=6(x-1)(x+6)\] whats the solution

OpenStudy (anonymous):

y=3(x-1)(x+6) y=3(-1-1)(-1+6) y=3(-2)(5) y=3(-10) y=-30 ??? is that right?? idk

OpenStudy (anonymous):

yes, so -1 cannot be correct, you want y to be 0

OpenStudy (anonymous):

@Jonask 0=6(-1-1)(-1+6) 0=6(-2)(5) 0=6(-10) 0=30 wait so what just happened there?

OpenStudy (anonymous):

@Rimplestorck so now C and D are out right??? so we need to consider between answer choices A and B?

hero (hero):

Wow, such confusion

OpenStudy (anonymous):

the y coordinates of x intercepts are 0, so yes, C and D are out

OpenStudy (anonymous):

haha yeah :/ @hero.. :P

OpenStudy (anonymous):

and how do we find the answer then?

OpenStudy (anonymous):

well you just need to find values of x for which 0=6(x−1)(x+6) is 0

OpenStudy (anonymous):

for example, let's take x-1 first

OpenStudy (anonymous):

what does x have to be for x-1 to be 0

OpenStudy (anonymous):

it has to be 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and for x+6 to be 0, x has to be?

OpenStudy (anonymous):

ohhh i see now.. it has to be -6 so my answer is B. (1,0) and (-6,0) ????

OpenStudy (anonymous):

yes indeed

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome!! thanks :) @Rimplestorck can u pls give @Jonask a medal for me?? i can only give one :/ thanks for the help!! :D

OpenStudy (anonymous):

i a;lready gave hero

OpenStudy (anonymous):

okay its all good then :) thanks tho :)